OpenStudy (anonymous):

Which number makes this equation true? v2 + 10v + 16 = (v + 8)(v + ?)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

v^2+10v+16 =v^2+8v+2v+16 =v(v+8)+2(v+8) =(v+8)(v+2)

OpenStudy (anonymous):

yw